What is the mass of a 1.82-L sample of a liquid that has a density of 0.921g/mL?
Answer #1

d= m/v

vol =1.82-L = 1820mL

density = 0.921g/mL

0.921g/mL= m/1820mL

m= 1676.22g

m= 1.646Kg